# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

### v2

@@ D.1 EIG Estimation
Suppose our model is given by $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t};[h_{t-1},x_{t}]\right)p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$}. Here, it will clearly be beneficial to directly use Equation˜1 for estimating the EIG, as here we directly have access to all the required terms, other than $p(\theta)$ which can be simply ignored as it is not a function of $x$ so does not affect optimization of the question. If the possible values for $y$ are enumerable and we can evaluate $p\left(y_{t};[h_{t-1},x_{t}]\right)$ in closed–form, we can directly calculate the exact EIG (up to a constant) without requiring any estimation at all:
@@
$\text{EIG}_{\theta}(x)-\mathrm{Const}=-\sum_{y}p\left(y_{t};[h_{t-1},x_{t}]\right)\mathrm{H}\left[p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right],$ (5)
where the entropy $\mathrm{H}\left[p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right]$ can be evaluated directly from the logits of the LLM, or if these are not available, estimated by sampling. If we cannot enumerate $y$ or evaluate $p\left(y_{t};[h_{t-1},x_{t}]\right)$, we can simply instead resort to Monte Carlo and use the estimator:
$\text{EIG}_{\theta}(x)-\mathrm{Const}\approx-\frac{1}{N}\sum_{n=1}^{N}\mathrm{H}\left[p(\theta;[c,x,y_{n}])\right]\quad\mathrm{where}\quad y_{n}\sim p\left(y_{t};[h_{t-1},x_{t}]\right).$ (6)
We provide an overview of how to do this in the LLM setting in Algorithm˜1.
D.2 Generating candidate hypotheses
@@
To generate candidate values of $\theta$ for the data–estimation method, we use the prompt in Section˜D.2.
@@
Is this person known for their contributions to science?
No.
Is this person known for their contributions to the arts?
Yes.
  "Vincent van Gogh": 93,
  "Salvador Dali": 44,
  "Frida Kahlo": 37,
  "Georgia O Keeffe": 10,
  "August Wilson": 8,
  "Auguste Rodin": 8
Figure 5: An example of the sample distribution generated using the prompt in Section˜D.2, conditioned on the two question/answer pairs at the top of this figure. At this stage of the game, we independently sample 200 hypotheses from the LLM and record their frequencies. Note that the distribution exhibits strong mode collapse, with most of the mass highly concentrated on just a few answers, which negatively impacts the performance of the data–estimation method. This summary is for diagnostic purposes: Algorithm˜1 operates on the probabilities of individual samples and never instantiates such an aggregated summary.
@@ Appendix E Generating candidate hypotheses for BED-LLM
Candidate hypotheses are generated jointly, rather than independently.
@@
We assume a uniform distribution over hypotheses
While one could in principle reweight candidates using importance sampling, in practice we choose to not rely on the model’s internal probabilities. Instead we approximate the prior as a uniform distribution over this union
@@
Appendix F Experiment details for 20 Questions
F.1 Problem sets
@@
To create the Animals problem set, we prompted OpenAI o3 (OpenAI, 2025, o3-2025-04-16) to generate a list of animals, using the prompt in Section˜F.1. The resulting list is shown in Figure˜8. Alternative names (after $|$ character) were manually added.
@@ F.2 Evaluation
Figure 9: Full plots for 20 Questions Experiments.
F.3 Algorithmic details
Using our sample–then–filter process (see Section˜3.3), we aim to sample at least $N=15$ hypotheses, repeating the cycle up to three times if needed (the exact number of hypotheses can be less than this as it may not be possible to generate sufficient valid hypotheses, especially in later experiment turns). The questioner generates $M=15$ candidate questions to test, $\mathcal{X}^{\text{cand}}$, using the “conditional generation” approach of Section˜3.1 when possible, but falling back on “unconditional generation” if insufficient candidate hypotheses have been generated.
Appendix G Experiment details for preference elicitation
G.1 Problems
To generate a set of ground-truth user profiles, we take a set of $200$ real user ratings from the MovieLens-100K dataset (Harper & Konstan, 2015), then use an “oracle” LLM (namely, OpenAI’s o3 model) to produce a paragraph of text that is consistent with each distinct set of ratings, using the prompt in Section˜G.3. As was the case for the 20 Questions problems, this problem set is never provided to the questioner and the set of allowed $\theta$ is not constrained.
@@ G.2 Evaluation
G.3 Algorithmic details
For BED-LLM and Entropy, we compare $M=8$ candidate questions at each turn and we aim to generate at least $N=5$ candidate hypotheses. We use the “unconstrained generation” approach of candidate question generation (see Section˜3.1) as the user profiles can be quite diffuse and we are only generating a small number of possible hypotheses that can be quite easy to split.

### v3

@@ Appendix D Data–estimation method
$p(\theta,y_{t};h_{t-1},x_{t})=p\!\left(y_{t};[h_{t-1},x_{t}]\right)p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right).$ (5)
@@ D.1 EIG estimation
Using the form of the EIG given by Equation˜1, we have
@@
$\mathrm{EIG}_{\theta}(x)=\underbrace{\mathrm{H}\!\left[p(\theta;h_{t-1},x)\right]}_{\text{marginal entropy of }\theta}\;-\;\underbrace{\mathbb{E}_{p\!\left(y_{t};[h_{t-1},x_{t}]\right)}\!\left[\mathrm{H}\left[p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right]\right]}_{\text{expected posterior entropy}}.$ (6)
Crucially, the marginal on $\theta$ implied by this pairing is obtained by integrating out $y$:
$p(\theta;h_{t-1},x)=\sum_{y_{t}}p\!\left(y_{t};[h_{t-1},x_{t}]\right)p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right),$ (7)
which is a mixture distribution whose components and weights both depend on $x$. Unlike in the prior–likelihood pairing, where the belief state $p_{f}(\theta;h_{t-1})$ is constructed independently of the candidate design, the marginal entropy $\mathrm{H}[p(\theta;h_{t-1},x)]$ here varies across candidate questions and cannot be dropped from the optimization. This is a concrete manifestation of the issue identified in Section˜3: the data–estimation pairing does not provide a design–independent belief state on $\theta$.
Discrete answer space.
When the possible values for $y$ are discrete (and finite) and we can evaluate $p\!\left(y_{t};[h_{t-1},x_{t}]\right)$ in closed form, both terms of Equation˜6 can be computed directly. The second term (expected posterior entropy) is:
$\mathbb{E}_{p\!\left(y_{t};[h_{t-1},x_{t}]\right)}\!\left[\mathrm{H}\!\left[p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right]\right]=\sum_{y_{t}}p\!\left(y_{t};[h_{t-1},x_{t}]\right)\mathrm{H}\!\left[p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)\right],$ (8)
where each posterior entropy $\mathrm{H}[p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)]$ can be evaluated from the LLM’s logits or, if these are unavailable, estimated by sampling. The first term (marginal entropy) requires evaluating $p(\theta;h_{t-1},x)$ as the mixture in Equation˜7, then computing its entropy. In practice, this entropy cannot be obtained in closed form and must be estimated. A natural approach is to draw samples $\theta_{m}\sim p(\theta;h_{t-1},x)$ (by first sampling $y_{m}\sim p\!\left(y_{t};[h_{t-1},x_{t}]\right)$, then $\theta_{m}\sim p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)[,y_{m}]$) and using a plug-in or nearest-neighbor entropy estimator, or evaluating $\log p(\theta_{m};h_{t-1},x)=\log\sum_{y}p\!\left(y_{t};[h_{t-1},x_{t}]\right)\,p(\theta_{m};[h_{t-1},x,y])$ exactly by enumerating over $y$.
Monte Carlo estimator.
When $y$ cannot be enumerated, both terms must be estimated by Monte Carlo. Drawing $N$ samples $y_{n}\sim p\!\left(y_{t};[h_{t-1},x_{t}]\right)$, the full estimator takes the form:
$\mathrm{EIG}_{\theta}(x;h_{t-1})\approx\;\mathrm{H}\!\left[\frac{1}{N}\sum_{n=1}^{N}p(\theta;[h_{t-1},x,y_{n}])\right]\;-\;\frac{1}{N}\sum_{n=1}^{N}\mathrm{H}\!\left[p(\theta;[h_{t-1},x,y_{n}])\right].$ (9)
Whereas Equation˜3 performs this computation over the answer space $\mathcal{Y}$ (which is small and enumerable for multiple-choice questions), Equation˜9 requires it over the hypothesis space $\Theta$. For the first term, evaluating the mixture entropy requires computing $p(\theta;[h_{t-1},x,y_{n}])$ for a set of sampled $\theta$ values across all $N$ mixture components — an $O(M\times N)$ cost in autoregressive probability evaluations over the full hypothesis text.
D.2 Implementation
We provide an overview of the Data–Estimation procedure in Algorithm˜1. In our experiments, the answer space $\mathcal{Y}$ is enumerable (multiple-choice options), so we use Equation˜8 for the second term. For the first term, we estimate the marginal entropy by sampling $\theta$ values from the mixture in Equation˜7 and evaluating $\log p(\theta;h_{t-1},x)$ by enumerating over $y$; that is, we use:
$\mathrm{H}\!\left[p(\theta;h_{t-1},x)\right]\approx-\frac{1}{M}\sum_{m=1}^{M}\log\!\left(\sum_{y}p\!\left(y_{t};[h_{t-1},x_{t}]\right)\,p(\theta_{m};[h_{t-1},x,y])\right),$ (10)
where $\theta_{m}$ are drawn from $p(\theta;h_{t-1},x)$. This requires $M\times|\mathcal{Y}|$ evaluations of $p(\theta_{m};[h_{t-1},x,y])$ per candidate question, each involving a full forward pass over the hypothesis text.
D.3 Generating candidate hypotheses
@@
To generate candidate values of $\theta$ for the data–estimation method, we use the prompt in Section˜D.3.
Conversation History ($\mathcal{D}$) Q: Is this person known for their contributions to science?   A: No. Q: Is this person known for their contributions to the arts?    A: Yes.
Sampled Hypothesis Distribution
Vincent van Gogh 93
Salvador Dali 44
Frida Kahlo 37
Georgia O’Keeffe 10
August Wilson 8
Auguste Rodin 8
Figure 5: An empirical hypothesis distribution generated by the LLM, conditioned on the conversation history (top). By independently sampling $N=200$ hypotheses, we observe significant mode collapse; the probability mass is heavily concentrated on a few prominent figures (e.g., Vincent van Gogh). This overconfidence negatively impacts the performance of the data-estimation baseline. Note that this aggregated view is purely diagnostic; Algorithm˜1 operates directly on the probabilities of individual samples.
@@
Section˜D.3 shows an example of in-context belief updating failing to respect the conversation history. The response to the first question indicates that the target is male; however, after the second question, the LLM generates two female candidate hypotheses.
Figure 7: An example of in-context belief tracking with GPT-4o-mini proposing hypotheses inconsistent with the history. Q1: Is the person male?
Answerer: Yes.
Q2: Is this person often associated with civil rights or social justice?
Answerer: Yes.
Sampled hypotheses: [’James Baldwin’, ’A. Philip Randolph’, ’Angela Davis’, ’Malcolm X’, ’W.E.B. Du Bois’, ’Desmond Tutu’, ’Cesar Chavez’, ’Rosa Parks’, ’Martin Luther King Jr.’, ’Langston Hughes’, ’John Lewis’, ’Frederick Douglass’, ’Nelson Mandela’, ’Thurgood Marshall’, ’Bayard Rustin’]
Hypotheses rejected with filtering: [’Angela Davis’, ’Rosa Parks’]
@@ Appendix E Generating candidate hypotheses for BED-LLM
Candidate hypotheses are generated jointly, rather than independently
@@
We define a uniform prior over hypotheses
While one could in principle reweight candidates using importance sampling, in practice we choose to not rely on the model’s internal probabilities. Instead we define the prior as a uniform distribution:
@@
Appendix F Additional experimental results
F.1 20 Questions and Preference Elicitation
Figure 9: Success rate on 20 Questions with Qwen2.5-72B, extended to 50 turns: mean $\pm$ standard error across 100 targets per dataset. BED-LLM remains the best-performing method at every turn. The performance gap to baselines narrows in later turns as the task saturates — most methods eventually solve the easier instances — but BED-LLM consistently retains its advantage.
Figure 10: Analysis of the empirical convergence of the Rao-Blackwellized EIG estimator using GPT-4o-mini. We plot EIG estimates for 5 candidate questions at turn 5 of the 20 Questions game with $\theta^{*}=$“Saoirse Ronan”, as a function of the number of samples in the hypothesis set $\Theta$.
Figure˜10 shows the EIG estimates for five candidate questions as the number of hypothesis samples $|\Theta|$ increases, at turn 5 of a representative 20 Questions game. The estimates converge rapidly with the number of samples, and more importantly, the ranking of candidate questions stabilizes well before the estimates themselves have fully converged. By $|\Theta|=10$ the relative ordering of all five candidates already matches the ordering observed at $|\Theta|=85$. This means that the question selection decision—which is what matters algorithmically, since we only need the $\arg\max$, not the absolute EIG value—is robust to using a small hypothesis set. This justifies BED-LLM’s use of a moderate sample budget without sacrificing decision quality.
Figure 11: Success rate on 20 Questions. Mean $\pm$ standard error across 100 targets. BED-LLM beats all other methods across all datasets and models evaluated.
Figure 12: Full plots for 20 Questions Experiments with different questioner and answerer.
Figure 13: Mean rating across 10 film recommendations: mean $\pm$ standard error across 200 users. BED-LLM beats all other methods across all datasets and models evaluated.
F.2 Wall-clock times
Method Runtime
Data–Estimation 1d 17h 13m 34s
ICL Beliefs 3h 10m 39s
Entropy 2h 46m 19s
Split 2h 30m 35s
BED-LLM 2h 28m 43s
Implicit Maximization 45m 10s
CoT 35m 28s
Prompt-Only 26m 29s
Table 3: Wall-clock runtimes for 20 Questions for Qwen2.5-72B on the entire Animals problem set.
Table˜3 shows that Data–Estimation is over an order of magnitude slower than all other methods: its full-sequence likelihood evaluations per candidate question dominate the cost. Among the remaining methods, BED-LLM’s runtime is comparable to—and even slightly lower than—Entropy and Split, despite computing the full EIG; this is because the additional conditional-entropy term in Equation˜3 reuses the same likelihood evaluations and adds negligible overhead. Prompt-Only and CoT are the cheapest methods, as they require only a single LLM generation call per turn with no hypothesis sampling or objective computation. Implicit Maximization falls in between: it samples hypotheses and candidate questions but replaces explicit EIG estimation with a single LLM selection call. ICL Beliefs is slightly more expensive than BED-LLM because the absence of filtering leads to lower-quality hypotheses, requiring a much larger number of questions before the game terminates.
F.3 Effect of sequentially updating the BED-LLM likelihood
As discussed in Section˜F.3, we choose to use a static likelihood, $p_{\mathrm{LLM}}\!\left(y_{t};\left[\theta,x_{t}\right]\right)$, in the model factorization for BED-LLM. Alternatively, we could condition the likelihood on the full interaction history, i.e. $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},\theta,x_{t+1}])$. However, this can lead to undesirable effects (e.g. context-induced calibration shifts). We provide results in Table˜4 for the 20 Questions game where we update the likelihood, and demonstrate that it leads to consistently worse performance than BED-LLM, supporting our rationale in Section˜F.3 to keep the likelihood static.
Success Rate (%)
Dataset Model BED-LLM + static likelihood BED-LLM + likelihood updating
Animals GPT-4o-mini 88 $\pm$ 3.3 85 $\pm$ 3.6
GPT-4o 93 $\pm$ 2.6 86 $\pm$ 3.5
Llama-3.1-8B 63 $\pm$ 4.9 67 $\pm$ 4.7
Llama-3.3-70B 79 $\pm$ 4.1 80 $\pm$ 4.0
Qwen2.5-72B 95 $\pm$ 2.2 89 $\pm$ 3.1
Celebrities GPT-4o-mini 72 $\pm$ 4.5 63 $\pm$ 4.9
GPT-4o 86 $\pm$ 3.5 83 $\pm$ 3.8
Llama-3.1-8B 58 $\pm$ 5.0 56 $\pm$ 5.0
Llama-3.3-70B 55 $\pm$ 5.0 44 $\pm$ 5.0
Qwen2.5-72B 84 $\pm$ 3.7 70 $\pm$ 4.6
Things GPT-4o-mini 49 $\pm$ 5.0 43 $\pm$ 5.0
GPT-4o 64 $\pm$ 4.8 57 $\pm$ 5.0
Llama-3.1-8B 26 $\pm$ 4.4 24 $\pm$ 4.3
Llama-3.3-70B 55 $\pm$ 5.0 54 $\pm$ 5.0
Qwen2.5-72B 62 $\pm$ 4.9 61 $\pm$ 4.9
Table 4: BED-LLM success rates across datasets and models. $\pm$ numbers show the standard error of the mean estimated using $\sqrt{p(1-p)/(n-1)}$ where $p$ is the success percentage and $n$ is the number of datapoints. This estimator is positively biased and thus conservative.
Appendix G Further description of methods and ablations
Prompt-Only: A direct prompting baseline where the model is asked to immediately generate the next question given the dialogue history, no candidate generation, and no belief modeling (Appendix˜G). This corresponds to the simplest and most common way users currently interact with LLM agents for question-asking tasks, and therefore serves as a natural “standard LLM prompting” baseline. After the LLM has generated its question, it waits for the user response and generates a new question based on the new history.
Figure 14: Prompt for the Prompt-Only method. Your task is to ask a series of questions to deduce the identity of the ’person’ if task == ’celebrities’ else ’animal’ if task == ’animals’ else ’entity’ that I’m thinking of with as few queries as possible.
Only ask questions that can be answered by ’yes’, ’no’ or ’maybe’.
Do not ask for hint. Make your question standalone with no linebreaker.
Output format: <question>Your question here?</question>
Now start asking a question.
CoT: A direct prompting baseline where the model is asked to reason and provide action (Appendix˜G) loosely following the ReAct framework (Yao et al., 2023b).
Figure 15: Prompt for the CoT method. Your task is to ask a series of questions to deduce the identity of the ’person’ if task == ’celebrities’ else ’animal’ if task == ’animals’ else ’entity’ that I’m thinking of with as few queries as possible.
Only ask questions that can be answered by ’yes’, ’no’ or ’maybe’.
Do not ask for hints. The actual question you ask must be standalone with no linebreakers.
Use the following format for every response:
Thought: <brief reasoning about what to ask next>
Action: ASK["<a single yes/no/maybe question with no linebreaks>"]
Now start by producing your first Thought and Action.
Split: Essentially a version of our method in which we cast previous approaches, such as Kobalczyk et al. (2025): uses the same candidate question generating method as BED-LLM, assumes a deterministic likelihood over hypotheses (i.e., each hypothesis deterministically predicts a single answer to each question), and Evaluates the question on the sampled hypotheses, and selects the question that maximally balances the hypotheses, i.e., splits the current hypothesis set into subsets whose sizes are as close to equal as possible.
Implicit Maximization (IM).
The Implicit Maximization baseline is a lightweight, reasoning-driven method inspired by the Tree-of-Thoughts (ToT, Yao et al., 2023a) framework. ToT methods explicitly expand a search tree of intermediate reasoning steps (“thoughts”) before selecting an action. Implicit Maximization provides a computationally efficient, collapsed variant of this idea:
  • Rather than explicitly branching over thoughts, the model is prompted to internally deliberate over several possible next questions and their consequences.
  • The branching and evaluation occur inside the model’s chain-of-thought, rather than through externally enumerated tree expansion.
  • The model then outputs the question it judges to be most informative, performing an amortized, ToT-style lookahead within a single LLM call.
Thus, IM captures the central intuition of Tree-of-Thoughts—reasoning over hypothetical futures—while avoiding the substantial computational cost of explicit tree search. It serves as a strong inference-time reasoning comparator to BED-LLM.
Appendix H Experiment details for 20 Questions
H.1 Problem sets
@@
To create the Animals problem set, we prompted OpenAI o3 (OpenAI, 2025, o3-2025-04-16) to generate a list of animals, using the prompt in Section˜H.1. The resulting list is shown in Figure˜17. Alternative names (after $|$ character) were manually added.
@@ H.2 Evaluation
H.3 Algorithmic details
Using our sample–then–filter process (see Section˜3.1), we aim to sample at least $N=15$ hypotheses, repeating the cycle up to three times if needed (the exact number of hypotheses can be less than this as it may not be possible to generate sufficient valid hypotheses, especially in later experiment turns). The questioner generates $M=15$ candidate questions to test, $\mathcal{X}^{\text{cand}}$, using the “conditional generation” approach of Section˜3.2 when possible, but falling back on “unconditional generation” if insufficient candidate hypotheses have been generated.
Appendix I Experiment details for preference elicitation
I.1 Problems
To generate a set of ground-truth user profiles, we take a set of $200$ real user ratings from the MovieLens-100K dataset (Harper & Konstan, 2015), then use an “oracle” LLM (namely, OpenAI’s o3 model) to produce a paragraph of text that is consistent with each distinct set of ratings, using the prompt in Section˜I.3. As was the case for the 20 Questions problems, this problem set is never provided to the questioner and the set of allowed $\theta$ is not constrained.
@@ I.2 Evaluation
I.3 Algorithmic details
For BED-LLM and Entropy, we compare $M=8$ candidate questions at each turn and we aim to generate at least $N=5$ candidate hypotheses. We use the “unconstrained generation” approach of candidate question generation (see Section˜3.2) as the user profiles can be quite diffuse and we are only generating a small number of possible hypotheses that can be quite easy to split.